What is the m&b of x+5y=5?
Thank you.

Answer by checkley77(12844)

Y=mX+b
x+5y=5
5Y=-X+5
Y=-X/5+5/5
Y=-X/5+1
m=-1/5 ANSWER.
b=1 ANSWER.